# Humanoid Locomotion and Manipulation: Current Progress and Challenges in Control, Planning, and Learning thanks: *co-corresponding authors

### v1

@@ VIII-B Building Humanoid Foundation Models
To leverage prior knowledge in FMs, another popular approach is to build a vision-language-action (VLA) model, as exemplified by RT-2 [357], OpenVLA [359], and Gato [379]. This approach treats robot data (i.e., observations and actions) as tokens in the language model’s vocabulary, allowing direct fine-tuning or co-training with existing VLMs. Unlike the previous approach, VLA outputs actions as tokens directly without relying on trainable low-level policies. The VLA model not only generates robot actions for diverse skills but also retains semantic reasoning abilities in the language and vision domain, enhancing its generalization capacity compared to models trained in single domains. For example, RT-2 [357] represents the actions as a string of numbers similar to the tokens from the pre-trained vision and language tokenizer of the base VLMs. However, representing actions with stringified numbers can be token-inefficient with high degrees-of-freedom robots such as humanoids. Despite the potential of this approach, no existing work has built VLA for humanoid robots. Recently, NVIDIA announced their Project GR00T initiative to develop general-purpose FMs for humanoid robots. The GR00T foundation model aims to leverage diverse data sources, from internet data and simulation data to real-robot data, for scalable training.
@@
A key research question in building RFMs is the design of effective algorithms and model architectures. Although most of today’s RFMs are based on autoregressive transformer models [378], their computational inefficiency over long sequences poses a significant challenge for both training and inference. This has driven exploration of alternative models that are both efficient and high-capacity, such as state-space models [380].
@@
The key to the success of training an RFM, especially in the case of VLA, heavily depends on the choice of input and output representations. Most RFMs take as input a combination of task descriptions in language, visual observations of the surrounding environment, and the history of robot states. Outputs typically consist of robot actions, derived either from a multi-task policy (the first approach) or a VLA model (the second approach). There are variations in the VLA model where its token outputs specify more than just actions. For example, Octo [381] and RDT-1B [382] use the token output for the diffusion denoising process. $\pi_{0}$ [375] maps a learned token to robot actions through a diffusion head, enabling high-frequency control (up to $50$ Hz for a bimanual manipulator with a wheelbase). GR-2 [383] predicts tokens that represent future images and actions; thus it functions as both a world model and a visuomotor policy. For humanoid robots, input and output are not yet well defined. For tasks involving rich physical interactions, the force feedback is as crucial as egocentric visual input. Given the unstable dynamics of humanoid robots, a more practical approach is to use end-effector and body poses as action outputs from the RFM and adopt additional low-level policies to ensure balance and safety via high-frequency feedback control. Incorporating robotic data as a new modality into state-of-the-art FMs would require significantly more data. Hence, designing effective input/output representation for humanoid loco-manipulation tasks still remains an open research question.
Conclusion: With rapid progress in FMs and humanoid robots, a plethora of research work on embodied intelligence in humanoid robots is anticipated. In the near term, we expect research to lean toward methods outlined in Sec. VIII-A as a more accessible way to leverage the capabilities of LLMs and VLMs. Meanwhile, we anticipate that the strategy of training a humanoid FM in Sec. VIII-B will become the mainstream in the long term as researchers develop a deeper understanding of incorporating additional modalities in FMs and with more humanoid robots deployed and data collected. For further reading of FMs for robotics, we recommend the survey in [9, 10]. Since most of the FMs are built with the Transformer backbones, please refer to [378] for a comprehensive mathematical foundation.
IX Future Challenges and Opportunities
IX-A Challenges in Numerical Optimization
Robotic planning and control techniques that are formalized as numerical optimization problems heavily rely on advances in applied discrete mathematics and optimization theory. These advancements address challenges such as non-convexity, numerical robustness, and real-time resolution performance. However, a plateau may have been reached in the transfer of these techniques to the field of robotics at large – humanoids specifically – despite exploiting the unique properties of robotic models. These properties can enhance the efficiency of optimization problem formulation and its resolution by tailoring them to specific applications. However, as evidenced by the formulation OCP, including WBC and MPC, inherent physical uncertainties can disrupt closed-loop performance. Extending these formulations to loco-manipulation primarily involves (i) augmenting models to incorporate loco-manipulated counterparts and (ii) refining contact models formulations to account for various interactions (e.g., impact [384], rolling, deforming). However, these extensions risk overcomplicating the problem, potentially hindering effective formalization, even if the resulting formulations are sparse.
Contact-explicit optimization formulations [6] are generally preferred due to their faster convergence and simplified formulation. However, they still suffer from the curse of dimensionality. Moreover, these formulations have the significant limitation of requiring the user to determine the contact mode sequence for the problem, which generally limits the ability to generate complex motions. Alternative contact-implicit formulations introduce complementarity conditions to eliminate the strict dependence on the contact mode sequence [160, 385, 165]. However, contact complementarity conditions are nonsmooth, introducing severe computational challenges. Generally, this is tackled via regularizing the complementarity problem, e.g.[168], which approximates the constraint with a continuous affine function. Even with this linearized approximation, contact-implicit approaches struggle to scale to the high dimensionality of humanoid robots due to excessive computation and numerical difficulty.
All these approaches, however, still have only local optimality guarantees—if the structure of the problem requires deviations from the local candidate contact conditions, a feasible solution, even one that is only locally optimal, may never be found. Additionally, they are almost always deterministic in nature, failing to capture the stochasticity in the state estimates and future contact events. Addressing this uncertainty and lack of global optimality has led to the combination of search techniques with traditional trajectory optimization. For example, Model Predictive Path Integral (MPPI) [212] samples a variety of random control signals and their resulting state trajectories to determine the best action to take. Alternative contacts and objectives can also be sampled with contact-implicit approaches to help avoid local minima and find the globally optimal solution to accomplish a task [386]. Although both of these approaches heavily leverage computational parallelization for expediency, parallelizing the underlying optimization algorithms explicitly designed for trajectory optimization is gaining increasing prevalence, both on the CPU [387] and GPU [388, 210], as discussed in Sec. V.
Despite the gains in these algorithms for considering the full system dynamics, the robustness of the mathematical solution when performing numerical optimization has been a concern due to the infeasibility of complex optimization problems, regardless of the computation speed improvement. As discussed in WBC Sec. VI, arbitrating techniques address infeasibility by (i) relaxing the hard constraint to the soft constraint by combining them in the cost with a weighted sum and (ii) prioritizing the constraints by achieving the important ones first. However, how to design a smart solver to automatically resolve this issue and provide numerical robustness is still an open question. In addition, weight-tuning in high-dimensional problems with complex objectives is nontrivial, highly task-dependent, and can lead to instability [263]. Researchers have made initial steps to apply auto-tuning techniques to streamline the tuning process in Optimal Control Problem (OCP) designs for humanoid robots [389, 390]. Until these are solved, assigning definitive but non-violating constraints and designing objective functions while maintaining global versatility still depend on expert knowledge.
IX-B Challenge: Lack of Benchmarks for Loco-manipulation
Humanoid loco-manipulation skills are at their infancy compared to other tasks such as humanoid locomotion [391] and tabletop manipulation [392]. Therefore, many simulators, such as Isaac Lab [216] or MuJoCo [215], lack large-scale and systematic benchmarks dedicated to humanoid loco-manipulation tasks. Developing well-designed tasks and evaluation metrics can significantly accelerate research in this field. Recently, HumanoidBench [280] and Mimicking-Bench [393] offer a list of loco-manipulation tasks. They provide a standardized benchmark for evaluating the performance of humanoid algorithms and help verify the reproducibility within MuJoCo and Isaac Gym. SMPLOlympics [394] presents a collection of sports as benchmarks for simulated humanoids and shows that leveraging human-motion demonstrations leads to human-like behaviors in these sports.
In addition to defining loco-manipulation tasks, the development of affordable and capable humanoid robots as standardized platforms for hardware evaluation can significantly accelerate the research. Research efforts on open-source humanoid hardware and software, such as Hector [2], the MIT Humanoid [181], the Berkeley Humanoid [395], and the Duke Humanoid [396] represent valuable contributions. In addition, an ecosystem in rapid prototyping and production of humanoid robots also accelerates hardware development, such as the Robotic Grand Factory [397]. The Robotic Grand Factory has incubated the Q-series humanoid robots [397] with the state-of-the-art capability of fast walking, multi-terrain adaptation, and explosive motions, enabling tasks such as organizing, storing, and reception.
IX-C Challenge: Data Scarcity
As discussed in Sec. VII-B, the four data sources present a trade-off between quality and availability. The lack of high-quality large-scale robot data becomes a bottleneck for robot skill learning. To solve the bottleneck, much effort has been put into data scaling. It is a heated debate in the community whether scaling is the road toward generalist humanoid robots.
We argue that generalization in loco-manipulation is achieved by including the motion of the manipulated objects, enhanced with a greater variety of sensing modalities instead of data quantity scaling. To develop truly versatile and adaptive humanoids, human data should also include that of the manipulated objects, cognitive actions (e.g., trust, compete, collaborate) paired with multimodal observations (e.g., whole-body haptic sensing, egocentric images), so that humanoids can learn the ‘how’ and ‘why’ from human data. However, instrumenting the environment and the manipulated objects with force and tactile sensing might be extremely complex. Therefore, research toward inferring force information from vision [399] and human-captured data [400] can be an intermediary solution to the problem of whole actors instrumentation. Recent work on collecting human multimodal observations [401, 402] and human kinetics [403] are a few examples that aim to bridge the gap between human animation and humanoid applications. Together with the rapid advancement in humanoid hardware, purposive learning with more informative human data will become the mainstream approach to achieve versatile and general-purpose humanoids.
IX-D Opportunities and Challenges in Foundation Models
@@
Integrating Foundation Models (FMs) into humanoid robots offers distinct opportunities and challenges. On the opportunity side, since the majority of data used to train FMs is collected by humans, the knowledge embedded in these models is inherently biased towards human-like embodiments. Consequently, humanoid robots could potentially utilize existing knowledge in FMs more effectively due to a smaller embodiment gap. This advantage extends beyond planning and control capabilities to include interactions with humans using natural modalities such as language and gestures. However, challenges arise from the humanoid form itself. First, the bipedal platform poses additional challenges in control and safety due to its inherent instability. Furthermore, it elevates expectations for naturalness in their movements and interactions, necessary to avoid the uncanny valley effect.
@@
Another challenge in applying FMs to humanoids arises from the high inference cost. Running large foundation models using only onboard computing is not feasible due to the limited power and computation, which hampers real-time hardware control. To address this challenge, several solutions have been proposed. One effective strategy involves adopting a decentralized hierarchy, where FMs operate over the cloud and provide only high-level decisions at a lower frequency, while another controller remains onboard and manages real-time task execution. However, inference delay and internet latency might impede the control performance. Another approach is to enhance the speed of the computing platform and the efficiency of FMs. For instance, NVIDIA introduced Jetson Thor, an on-board computing platform designed for humanoid robots. Google proposed SARA-RT [404], which accelerates the model speed without compromising its quality.
@@
The training of FM is also resource and time consuming. For example, training the LLaMA model took $34$ days on $992$ NVIDIA A100-80B GPUs [405], which incurs high cost, high energy consumption, as well as carbon dioxide emission. As FMs scale up further, the training cost would increase further. A promising approach to maintain a reasonable cost for training robotics FMs is to leverage parameter-efficient fine-tuning techniques. For example, OpenVLA leverages the LoRA technique to fine-tune an FM with a robotics dataset [359].
@@ X Conclusion
Humanoid robotics is advancing at an unprecedented pace, as seen in recent groundbreaking innovations from both industry and academia. Although humanoid robots still face significant technical challenges, their agility, safety, reliability, and versatility have improved significantly through both model-based and learning-based methods. More opportunities are emerging through the exploration of new paradigms. Physically, advanced observers (vision and whole-body tactile sensing) and estimators are emerging for contact-rich whole-body loco-manipulation. Cognitively, foundation models grounded on humanoid robots have great potential to unlock the ability of open-world understanding and the development of generalized intelligent agents. More importantly, how to seamlessly integrate multimodal sensing and foundation models with the existing planning and control frameworks presents promising and challenging research questions. In the foreseeable future, the cost of humanoid robots will continue to decrease, making them more accessible; their physical capability (hardware intelligence) and cognitive intelligence will significantly advance. We look forward to humanoid robots that are responsive, purposeful, and fully capable of human-like loco-manipulation tasks in the upcoming decade.

### v2

@@ VIII-B2 End-to-end Vision-Language-Action model
To leverage prior knowledge in FMs, another popular approach is to build an end-to-end VLA model, as exemplified by RT-2 [385], OpenVLA [387], and Gato [408]. This approach treats robot data (i.e., observations and actions) as tokens in the language model’s vocabulary, allowing direct fine-tuning or co-training with existing VLMs. Unlike the previous approach, VLA outputs actions as tokens directly without relying on trainable low-level policies. The VLA model not only generates robot actions for diverse skills, but also retains semantic reasoning abilities in the language and vision domain, enhancing its generalization capacity compared to models trained in single domains. For example, RT-2 [385] represents the actions as a string of numbers similar to the tokens from the pre-trained vision and language tokenizer of the base VLMs. However, representing actions with stringified numbers can be token-inefficient with high degrees-of-freedom robots such as humanoids.
@@
A key research question in building RFMs is the design of effective algorithms and model architectures. Although most of today’s RFMs are based on autoregressive transformer models [405], their computational inefficiency over long sequences poses a significant challenge for both training and inference. This has driven the exploration of alternative models that are both efficient and high-capacity, such as state-space models [409].
@@
The key to the success of training an RFM, especially in the case of VLA, heavily depends on the choice of input and output representations. Most RFMs take as input a combination of task descriptions in language, visual observations of the surrounding environment, and the history of robot states. Outputs typically consist of robot actions, which are derived from either a multi-task policy (the first approach) or an end-to-end VLA model (the second approach). There are variations in the VLA model where its token outputs specify more than just actions. For example, Octo [410] and RDT-1B [411] use the token output for a diffusion denoising process. $\pi_{0}$ [406] maps a learned token to robot actions through a diffusion head, enabling high-frequency control (up to $50$ Hz for a bimanual manipulator with a wheelbase). GR-2 [412] predicts tokens that represent future images and actions; thus it functions as both a world model and a visuomotor policy. For humanoid robots, input and output are not yet well defined. For tasks involving rich physical interactions, the force feedback is as crucial as egocentric visual input. Given the unstable dynamics of humanoid robots, a more practical approach is to use end-effector and body poses as action outputs from the RFM and adopt additional low-level policies to ensure balance and safety via high-frequency feedback control. Incorporating robotic data as a new modality into state-of-the-art FMs would require significantly more data. Hence, designing effective input/output representation for humanoid loco-manipulation tasks still remains an open research question.
VIII-C Opportunities and Challenges in Foundation Models
@@
Integrating Foundation Models (FMs) into humanoid robots offers distinct opportunities and challenges. On the opportunity side, since the majority of data used to train FMs is collected by humans, the knowledge embedded in these models is inherently biased towards human-like embodiments. Consequently, humanoid robots could potentially utilize existing knowledge in FMs more effectively due to a smaller embodiment gap. This advantage extends beyond planning and control capabilities to include interactions with humans using natural modalities such as language and gestures.
@@
A major challenge in applying FMs to humanoids arises from the high inference cost. Running large foundation models using only onboard computing is not feasible due to the limited power and computation, which hampers real-time hardware control. To address this challenge, several solutions have been proposed. One effective strategy involves adopting a decentralized hierarchy [388], where FMs operate over the cloud and provide only high-level decisions at a lower frequency, while another controller remains onboard and manages real-time task execution. However, the inference delay and internet latency may impede the control performance. Another approach is to enhance the speed of the computing platform and the efficiency of FMs. For instance, NVIDIA introduced Jetson Thor, an onboard computing platform designed for humanoid robots. Google proposed SARA-RT [413], which accelerates the model speed without compromising its quality.
@@
The training of FM is also resource and time consuming. For example, training the LLaMA model took $34$ days on $992$ NVIDIA A100-80B GPUs [414], which incurs high cost, high energy consumption, as well as carbon dioxide emission. As FMs continue to scale up, the training cost increases further. A promising approach to maintaining a reasonable cost for training robotics FMs is to leverage parameter-efficient fine-tuning techniques. For example, OpenVLA leverages the LoRA technique to fine-tune an FM with a robotics dataset [387].
@@
Beyond computational limitations, FMs face additional challenges when deployed in humanoid systems. Safety constraints are particularly concerning, as models trained on internet data may generate unsafe actions when controlling physical robots that are inherently unstable. Establishing a thorough evaluation procedure to assess the safety level of humanoid robots will be essential. Moreover, ethical considerations arise from biases present in human-generated training data, which can lead to unfair or discriminatory robot behaviors across different user demographics. These biases may be amplified in embodied systems that interact directly with humans in diverse settings. Federated learning approaches could help address some of these ethical concerns by enabling model training across distributed devices while preserving data privacy and potentially reducing centralized data biases [418].
Key Takeaways: With rapid progress in FMs and humanoid robots, a plethora of research studies on humanoid embodied intelligence is anticipated. The first method outlined in Sec. VIII-A provides a more accessible way to leverage VLMs and, thus, has been the popular approach taken by researchers so far, but VLMs as is lack a deep understanding of robot actions. In the near future, we expect increased research efforts devoted to VLA models described in Sec. VIII-B to address this issue. It takes great effort to train RFMs from (i) a large amount of real robot data (e.g., Gemini-Robotics [388, 389]) or (ii) a combination of robot data with synthetic data from simulation or video generation models (e.g., GR00T N1 [377]). With the development of RFMs, we expect to have FMs that can well understand robot action and the physical world, enabling rapid deployment of robotic controllers achieved with little to no additional robot data. For further reading of FMs for robotics, we recommend the survey in [11, 12]. Since most of the FMs are built with the Transformer backbones, please refer to [405] for a comprehensive mathematical foundation.
IX Additional Discussions
IX-A Model-based Methods Versus Learning-based Methods
Based on the studies in this survey, we provide a clear comparison between model-based and learning-based methods in Table. VI. In this table, we compare the performance in several aspects: the ability to incorporate multi-modal sensor input, the control robustness, the motion accuracy, the real-time feasibility, the versatility, and the generalizability of each method. Additionally, we evaluate the state-of-the-art capability in several humanoid skills, including locomotion, loco-manipulation, and whole-body multi-contact control. We list representative papers that achieved these skills as supporting evidence for this evaluation. Note that many algorithm performances reported are anecdotal and highly depend on their implementation. Therefore, obtaining a concrete numerical comparison is challenging. This is mentioned in Sec. VII-G challenges in lack of appropriate benchmark.
A key takeaway from this comparison is that learning-based methods exhibit superior robustness via RL and versatility via IL, while model-based approaches offer advantages in motion accuracy. The emerging paradigm of VLA holds promise for enhanced generalizability; however, its effectiveness in mastering humanoid skills has yet to be convincingly demonstrated.
TABLE VI: Comparison between model-based and learning-based methods
Model-based Learning-based
MPC, WBC RL IL VLA
Performance measure
Flexibility to incorporate
multi-modal sensor
low medium high very high
Robustness high very high low very low
Motion accuracy very high medium medium very low
Real-time feasibility high high high low
Versatility high medium high high
Generalizability medium low low high
Achieved performance of humanoid skills
Locomotion
very high
[6, 30]
very high
[40, 37]
medium
[283]
NA
Loco-manipulation
(hand and foot contact)
high
[2, 218]
medium
[41, 281]
low
[268]
NA
Whole-body
multi-contact
medium
[140]
high
[419]
NA NA
IX-B Hardware Limitations for Humanoid Robots
Advancements in control, planning, and learning for loco-manipulation necessitate robot hardware that has high-power-density actuators, strong yet lightweight limbs, and accurate sensing. These requirements are often limitations to hardware performance, which drives hardware designers to push the boundary of humanoid physical capabilities.
The traditional approach to increase torque capacity is through high gear reduction [420] and the inclusion of elastic elements to be impact-resilient [28]. However, the last decade has witnessed advancements in electric motor design such as Quasi-Direct Drive (QDD) actuators (less than 10:1 gear ratio) achieving dynamic and accurate motions [368, 2]. Among the many advantages QDD actuators offer for achieving dynamic motions are their backdrivability and high force control bandwidth, due to reduced actuator backlash and friction [179]. However, achieving high torque requires a high current, which in turn causes strain on power electronics and overheating. Solving this problem involves manufacturing with materials to enhance heat dissipation [420] and continuing to improve controller performance for energy efficiency [43].
Another key element of hardware design is the robot’s inertial properties. In building robots capable of gymnastic-style maneuvers, many designs place mass away from end-effectors [179, 2, 46]. The MIT Humanoid, for example, contains around $75$% of its mass in the torso [179]. This is often achieved through belt drives or parallel mechanisms to remotely actuate joints such as the ankle or knee [2, 46]. While this significantly reduces rotational inertia and the torque demand for agile maneuvers, there are a number of notable drawbacks. Four-bar transmissions often reduce joint range of motion, while belt drives may require additional modeling and maintenance [420]. Specialized mechanisms for remote actuation that allow a full range of motion can potentially address this problem [421].
Battery technology remains one of the primary hardware limitations in the development of sustained, autonomous humanoid robots. Limited battery capacity restricts both operational time and the feasibility of executing high-power, full-body motions in untethered scenarios. One promising direction is the development of more compact and integrated battery designs. For example, embedding battery cells directly into the structural components of the robot chassis can increase effective energy capacity without compromising weight distribution or form factor. In parallel, solid-state batteries have become commercially available with enhanced energy density and safety profiles [422]. This has increased the operation times of humanoid robots, such as Apptronik’s Apollo (4hrs) [423] and Unitree’s G1 (2hrs) [424]. Nevertheless, the continued reliance on tethered power in laboratory settings [5, 1, 46, 4] indicates that battery technology remains a limitation.
IX-C The Future Applications and Challenges of Humanoid Robots in the Society
Humanoid robots with loco-manipulation capabilities will become an essential part of our lives and a pillar of the future economy, similar to how smartphones and autonomous driving have transformed our lives.
Given a sufficient amount of time, humanoid robots will be increasingly autonomous and capable of performing general human-level tasks across a wide range of domains. The components discussed in this paper—sensing, planning, control, and learning—coordinate in real time to execute complex, goal-directed behaviors. This integration enables humanoid robots to take on tedious, hazardous, and physically demanding tasks.
  • In outdoor environments, they could serve as first responders during national crises, providing medical care and emergency support; monitor controlled settings in civil and mechanical infrastructure; deliver and sort building materials at construction sites; explore subterranean or collapsed structures with harsh and unpredictable terrain; and assist in agricultural tasks such as planting crops.
  • In factories and warehouses, humanoid robots have the potential to address the growing demand for flexible manufacturing by performing tasks such as delivery, assembly, and inspection, helping to offset the impact of a declining industrial workforce.
  • In human-centered settings such as households and hospitals, they have the potential to carry laundry, load dishwashers, and move patients between beds, supporting the needs of an aging population and improving their quality of life.
However, significant challenges remain in integrating humanoid robots into all aspects of society. To be fully accepted, these robots must be physically and mentally reliable, highlighting the core technical challenges of achieving robust and safe mobility and manipulation. Their intentions must also be well predicted and communicated with humans, which calls for advancements in effective human-robot interaction. Beyond the technical hurdles, economic and ethical concerns also pose major obstacles. Humanoid robots are still far from being commercially viable due to high costs, and the absence of legal and regulatory policies further hinders their appearance in our daily lives.
@@ X Conclusion
Humanoid robotics is advancing at an unprecedented pace, as seen in recent groundbreaking innovations from both industry and academia. In this survey, we go over foundational and state-of-the-art methods for humanoid loco-manipulation, from both model-based and learning-based perspectives. Although humanoid robots still face significant technical challenges, their agility, safety, reliability, and versatility have been improved significantly. More opportunities are emerging through the exploration of new paradigms. Physically, advanced observers (vision and whole-body tactile sensing) and estimators are emerging for contact-rich whole-body loco-manipulation. Cognitively, foundation models grounded on humanoid robots have great potential to unlock the ability of open-world understanding and the development of generalized intelligent agents. More importantly, an integration of foundation models with planning and control policies presents promising capabilities. In the foreseeable future, the cost of humanoid robots will continue to decrease, making them more accessible; their physical capability (hardware intelligence) and cognitive intelligence will significantly advance. We look forward to humanoid robots that are responsive, purposeful, and fully capable of human-like loco-manipulation tasks in the upcoming decade.